# Graph Condensation: A Survey

### v1

@@ 5 Condensed Graph Generation
Different from traditional graph reduction methods that trim the original graph, GC focuses on creating an entirely new condensed graph. An effective graph generation method not only accelerates optimization convergence but also enhances overall performance. In this section, we discuss the condensed graph generation methods from two aspects: attribute initialization and structure construction, as shown in Fig. 7.
Figure 7: Condensed graph generation methods.
5.1 Attribute Initialization
@@
Diverse initializations of the attribute matrix $\mathbf{X}^{\prime}$ can lead to varying convergence performance. To expedite the optimization process for the attribute matrix, researchers have investigated a range of initialization methods as substitutes for the conventional random initialization. For instance, GCond Jin et al. (2022b) initiates attribute matrix by randomly selecting original nodes with corresponding labels. SFGC Zheng et al. (2023c) employs the coreset method to select the most representative nodes for initialization. Besides the node selection, Mcond Gao et al. (2024) and CTRL Anonymous (2023) demonstrate that employing the synthetic embeddings can further enhance optimization performance. Accordingly, these methods first cluster original graph nodes based on the pre-defined compression rate, and then utilize the cluster centroids as the initialization of attribute matrix.
5.2 Structure Construction
@@
The adjacency matrix $\mathbf{A}^{\prime}$ of the condensed graph reflects relationships among condensed nodes and preserves the topological information of the original graph. The construction method of adjacency matrix affects its characteristics, such as homophily, spectral properties, and sparsity, which in turn impact overall performance. Current studies employ three kinds of strategies for generating adjacency matrices: generative model, parameterization, and predefined structure.
@@
Generative model. A straightforward generation method involves employing the generative model to construct adjacency matrix explicitly. For example, SGDD Yang et al. (2023a) utilizes noise, along with the attributes and labels of the condensed graph, as inputs to the generative model for creating the adjacency matrix. Although this method offers greater flexibility in controlling the generated structures, it suffers from the substantial modeling complexity: the number of parameters in $\mathbf{A}^{\prime}$ grows quadratically with the increase of condensed nodes.
@@
Parameterization. To mitigate modeling complexity of graph structure, several GC methods propose to parameterize the adjacency matrix. GCond Jin et al. (2022b) constructs the adjacency matrix based on homophily and models the adjacency matrix by node similarities. This parameterization strategy is designed to investigate the correlations between pairs of condensed nodes. KIDD Xu et al. (2023) directly optimize the low-rank decomposed matrices of the adjacency matrix to speed up the convergence. To preserve the spectral characteristics, GCEM Liu et al. (2023a) decomposes the adjacency matrix into eigenbasis and constructs the condensed structure using the eigenvectors matched with the most significant eigenvectors of the original graph.
@@
Pre-defined structure. To further simplify the modeling of condensed graph, several studies leverage predefined structure. For instance, MSGC Gao and Wu (2023) pre-defines multiple sparse adjacency matrices for condensed nodes. By this means, MSGC reduces the complexity of the adjacency matrix while capturing a more diverse range of structural information compared to the single dense matrix. Moreover, several GC methods have incorporated the structure-free strategy Jin et al. (2022b); Zheng et al. (2023c); Liu et al. (2022, 2023b, 2023c) by employing a fixed identity matrix as the adjacency matrix. This strategy leads to a significant simplification of the computational requirements for adjacency matrix.
5.2.1 Condensed Graph Sparsification
@@
The structure construction methods typically yield dense adjacency matrices with continuous values. These dense matrices often contain small but storage-intensive values that have minimal impact on GNN aggregation performance. Moreover, certain application scenarios, such as molecular graphs, require the binary and discrete adjacency matrices for downstream tasks. Hence, it becomes essential to sparsify the generated adjacency matrices, and numerous sparsification methods have been proposed to tackle these challenges. For instance, GCond Jin et al. (2022b) utilizes the threshold to remove the entries with small values and justifies that suitable choices of threshold do not degrade performance but increase the scalability of condensed graph. SGDD Yang et al. (2023a) employs the L2 norm of the adjacency matrix as a regularization term for sparsity control. To construct binary and discrete adjacency matrices, DosCond Jin et al. (2022a) and KIDD Xu et al. (2023) model each pair of nodes as an independent Bernoulli variable. Then, the Gumbel-Max reparametrization trick Maddison et al. (2016) is utilized to address the non-differentiability challenges in the optimization procedure.
6 Applications
@@
Due to the efficiency in training GNN models and capacity to conserve storage costs, GC contains vast application prospects and is increasingly being utilized across a variety of fields. In this section, we enumerate the diverse applications of GC, including hyperparameter/neural architecture search, graph continual learning, federated learning, inference acceleration and recommender systems.
6.1 Hyper-Parameter/Neural Architecture Search
@@
The hyper-parameter/neural architecture search Gao et al. (2021) involves systematic experimentation with a range of hyper-parameters or architectures to identify the optimal combination for downstream task. This process typically requires the training of multiple models with varying hyper-parameters/architectures on the same dataset, incurring substantial computational costs. To mitigate the computational burden, HCDC Ding et al. (2022) introduces GC in this problem and additionally generates the synthetic validation data by matching the hyper-parameter gradients. This ensures that the performance ranking of hyper-parameters/architectures on condensed datasets is consistent with that on the original datasets. By integrating GC, HCDC significantly reduces the time required for the search process.
6.2 Graph Continual Learning
@@
Graph continual learning Febrinanto et al. (2023) is the approach that focuses on the continuous acquisition of knowledge from the dynamic and evolving graph data. In this realm, a critical challenge is the catastrophic forgetting problem, where a model’s performance on previously learned data deteriorates significantly when trained on new data. To mitigate this issue, CaT Liu et al. (2023b) and PUMA Liu et al. (2023c) incorporate GC into the process of graph continual learning, leveraging the condensed graphs to preserve historical information within a memory bank. To maintain a balance between historical and newly incoming data, CaT strategically trains the model exclusively on the condensed graph. This method not only addresses the challenge of data imbalance in graph continual learning but also significantly enhances learning efficiency.
6.3 Federated Learning

### v2

@@ 5 Condensed Graph Generation
Different from traditional graph reduction methods that trim the original graph, GC focuses on creating an entirely new condensed graph. An effective graph generation method not only accelerates optimization convergence but also enhances overall performance. In this section, we discuss the condensed graph generation methods from two aspects: feature and structure generation.
5.1 Feature Generation
@@
The feature matrix $\mathbf{X}^{\prime}$ dominates parameter quantity in the condensed graph and the most of methods update it by gradients. As a result, the feature matrix is significantly influenced by diverse initializations, potentially leading to varied convergence speeds. To expedite the optimization process for the feature matrix, researchers have investigated a range of initialization methods as substitutes for conventional random initialization. For instance, GCond [25] initiates the feature matrix by randomly selecting original nodes with corresponding labels. SFGC [28] leverages the coreset method [89] and chooses the representative nodes for initialization. On the other hand, MCond [26] and CTRL [51] demonstrate that employing the representative embeddings can accelerate the convergence of optimization. Accordingly, these methods first cluster original graph nodes based on the pre-defined compression rate, and then utilize the cluster centroids as the initialization of the feature matrix.
In addition to gradient-based optimization, which relies on different initializations, CGC [66] proposes generating the feature matrix through a closed-form solution. This approach eliminates the need for iterative updates of the feature matrix, thus ensuring a more precise and efficient feature generation process.
5.2 Structure Generation
@@
The adjacency matrix $\mathbf{A}^{\prime}$ of the condensed graph reflects relationships among condensed nodes and preserves the topological information of the original graph. The construction method of the adjacency matrix affects its characteristics, such as homophily, spectral properties, and sparsity, which in turn impact overall performance. As shown in Table III, current studies employ three kinds of strategies for generating adjacency matrices: generative model, parameterization, and predefined structure.
@@
Generative model. A straightforward generation method involves employing the generative model to construct an adjacency matrix explicitly. For example, GCond [25] constructs the adjacency matrix based on homophily and generates the adjacency matrix by node similarities. The generative model is designed to investigate the correlations between pairs of condensed nodes. SGDD [30] utilizes noise, along with the features and labels of the condensed graph, as inputs to the generative model for creating the adjacency matrix. DisCo [67] transfers the link prediction model of the original graph to the condensed nodes, generating the corresponding edges for the condensed graph. Although this method offers greater flexibility in controlling the generated structures, it suffers from substantial modeling complexity: the number of parameters in $\mathbf{A}^{\prime}$ grows quadratically with the increase of condensed nodes.
@@
Parameterization. To mitigate the modeling complexity of graph structure, several GC methods propose to parameterize the adjacency matrix. GCSR [60] explores the self-expressive property among condensed nodes and further enhances the condensed graph by transferring the inter-class correlations from the original graph structure. Besides, KiDD [29] directly optimizes the low-rank decomposed matrices of the adjacency matrix to speed up the convergence. To preserve the spectral characteristics, GCEM [64] decomposes the adjacency matrix into an eigenbasis and constructs the condensed structure using the eigenvectors matched with the most significant eigenvectors of the original graph.
@@
Pre-defined structure. To further simplify the modeling of condensed graphs, several studies leverage pre-defined structures for the condensed graph. For instance, MSGC [48] pre-defines multiple sparse adjacency matrices for condensed nodes. By this means, MSGC reduces the complexity of the adjacency matrix while capturing a more diverse range of structural information compared to the single dense matrix. Moreover, several GC methods have incorporated the structure-free strategy [25, 28, 63, 27, 65] by employing a fixed identity matrix as the adjacency matrix. This strategy leads to a significant simplification of the computational requirements for the adjacency matrix.
TABLE IV: The accuracy (%) comparison between GC methods under different compression rates. “OOM” means out-of-memory. The highest accuracies are highlighted in bold, and runners-up are underlined. “Original graph” denotes the GNN performance which is trained on the original graph.
Dataset $r$ GCond GCDM GC-SNTK SFGC
Original
graph
Cora 1.30% 79.8±1.3 69.4±1.3 81.7±0.7 80.1±0.4 81.2±0.2
2.60% 80.1±0.6 77.2±0.4 81.5±0.7 81.7±0.5
5.20% 79.3±0.3 79.4±0.1 81.3±0.2 81.6±0.8
Citeseer 0.90% 70.5±1.2 62.0±0.1 66.4±1.0 71.4±0.5 71.7±0.1
1.80% 70.6±0.9 69.5±1.1 68.4±1.1 72.4±0.4
3.60% 69.8±1.4 69.8±0.2 69.8±0.8 70.6±0.7
Ogbn-arxiv 0.05% 59.2±1.1 59.3±0.3 64.4±0.2 65.5±0.7 71.4±0.1
0.25% 63.2±0.3 59.6±0.4 65.1±0.8 66.1±0.4
0.50% 64.0±0.4 62.4±0.1 65.4±0.5 66.8±0.4
Flickr 0.10% 46.5±0.4 46.1±0.1 46.7±0.1 46.6±0.2 47.2±0.1
0.50% 47.1±0.1 46.8±0.1 46.8±0.1 47.0±0.1
1.00% 47.1±0.1 46.7±0.1 46.5±0.2 47.1±0.1
Reddit 0.05% 88.0±1.8 89.3±0.1 OOM 89.7±0.2 93.9±0.0
0.10% 89.6±0.7 89.7±0.2 OOM 90.0±0.3
0.20% 90.1±0.5 90.2±0.4 OOM 89.9±0.4
@@
Condensed graph sparsification. The structure construction methods typically yield dense adjacency matrices with continuous values. These dense matrices often contain small but storage-intensive values that have minimal impact on GNN aggregation performance. Moreover, certain application scenarios, such as molecular graphs, require binary and discrete adjacency matrices for downstream tasks. Hence, it becomes essential to sparsify the generated adjacency matrices, and numerous sparsification methods have been proposed to tackle these challenges. For instance, GCond [25] utilizes the threshold to remove the entries with small values and justifies that suitable choices of threshold do not degrade performance but increase the scalability of the condensed graph. SGDD [30] employs the L2 norm of the adjacency matrix as a regularization term for sparsity control. To construct binary and discrete adjacency matrices, DosCond [52] and KiDD [29] model each pair of nodes as an independent Bernoulli variable. Then, the Gumbel-Max reparametrization trick [90] is utilized to address the non-differentiability challenges in the optimization procedure.
6 Empirical Studies
Considering the diversity of optimization strategies and GC criteria, it is crucial to understand the characteristics of different GC methods and design appropriate modules for enhancing specific task performance. Although this is not a paper centered on experiments and analysis, in this section we empirically analyze existing GC methods with diverse optimization strategies in terms of the five proposed criteria to provide the guidelines for both researchers and practitioners.
TABLE V: The generalizability comparison of GC methods. GC-SNTK is out-of-memory on Reddit. “Avg.” indicates the average value. The highest accuracies are highlighted in bold, and runners-up are underlined.
Dataset Method SGC GCN SAGE APPNP Cheby GAT Avg.
Cora ($r$ = 2.60%) GCond 79.3 80.1 78.2 78.5 76.0 66.2 76.4
GCDM 79.0 77.2 54.4 77.5 72.7 60.4 70.2
GC-SNTK 79.4 81.5 79.1 77.4 73.5 71.9 77.1
SFGC 79.1 81.7 81.9 78.8 79.0 80.8 80.2
Citeseer ($r$ = 1.80%) GCond 70.3 70.6 66.2 69.6 68.3 55.4 66.7
GCDM 72.0 69.5 67.1 69.3 68.7 61.7 68.1
GC-SNTK 71.6 68.4 66.5 69.4 67.6 61.9 67.6
SFGC 71.8 72.4 71.7 70.5 71.8 72.1 71.7
Ogbn-arxiv ($r$ = 0.25%) GCond 63.7 63.2 62.6 63.4 54.9 60.0 61.3
GCDM 61.2 59.6 61.1 62.8 55.4 61.2 60.2
GC-SNTK 62.7 65.1 62.9 62.6 55.1 61.8 61.7
SFGC 64.8 66.1 64.8 63.9 60.7 65.7 64.3
Flickr ($r$ = 0.50%) GCond 46.1 47.1 46.2 45.9 42.8 40.1 44.7
GCDM 44.3 46.8 45.8 45.2 41.8 41.9 44.3
GC-SNTK 45.7 46.8 45.9 45.3 41.3 41.4 44.4
SFGC 47.0 47.0 42.5 40.7 45.4 45.3 44.7
Reddit ($r$ = 0.10%) GCond 89.6 89.6 89.1 87.8 75.5 60.2 82.0
GCDM 88.0 89.7 89.3 88.9 74.9 69.3 83.4
SFGC 89.5 90.0 90.3 88.3 82.8 87.1 88.0
6.1 Evaluation Metric
Each criterion for the GC method is evaluated by a specific metric.
Effectiveness is assessed by comparing the accuracy of GNNs trained on condensed graphs at varying compression rates to those trained on the original graph. Methods that closely match the accuracy of the original setup are considered more effective.
Generalization is measured by comparing the average performance of models trained on condensed graphs across various GNN architectures and a range of downstream tasks, such as node classification, link prediction, anomaly detection, and temporal prediction. For a specific task $t$, the average performance of $n$ GNN architectures is defined as:
$p_{t}=\frac{1}{n}\sum_{i=1}^{n}acc_{i},$ (10)
where $acc_{i}$ represents the accuracy of the $i^{th}$ GNN architecture. A higher value indicates superior generalization ability for task $t$.
Efficiency is evaluated by the total time required for the condensation procedure, with shorter times indicating better practical usability.
Fairness is assessed by comparing the bias performance of GNNs trained on condensed graphs to those trained on the original graph. For the sensitive feature $s\in\{0,1\}$, it is desired that GNN predictions remain independent of the sensitive feature [49, 91]. With binary labels denoted as $y\in\{0,1\}$, and model predictions as $\hat{y}\in\{0,1\}$, model bias is quantified by demographic parity ($\Delta_{DP}$) [92] and equal opportunity ($\Delta_{EO}$) [93]:
$\small\Delta_{DP}=\left|P(\hat{y}=1|s=0)-P(\hat{y}=1|s=1)\right|,$ (11)
$\small\Delta_{EO}=\left|P(\hat{y}=1|{y}=1,s=0)-P(\hat{y}=1|{y}=1,s=1)\right|.$ (12)
GC methods that effectively mitigate the bias of GNNs trained on the original graph are considered fairer.
Robustness is evaluated by the accuracy of GNNs trained on condensed graphs derived from original graphs with varying noise levels. Methods that achieve higher accuracy are considered to preserve core information more effectively and demonstrate greater robustness against noise in the original graph.
6.2 Experimental Settings
Methods. We compare four representative GC methods with different optimization strategies: (1) gradient matching-based method: GCond [25]; (2) trajectory matching-based method: SFGC [28]; (3) KRR-based method: GC-SNTK [61]; (4) distribution matching-based method: GCDM [63].
Datasets. Given the diversity of evaluation criteria, GC methods are evaluated across various datasets with distinct characteristics, including three transductive datasets (Cora, Citeseer [94] and Ogbn-arxiv [95]), two inductive datasets (Flickr and Reddit [15]) and a dataset with sensitive feature (Pokec-n [91]).
Implementations. Following GCond [25], we evaluate three compression rates ($r$) for each dataset to evaluate the effectiveness. In the transductive setting, $N$ represents the original graph size, while in the inductive setting, $N$ indicates the sub-graph size observed in the training stage. Two-layer GNNs with 256 hidden units are used for evaluation.
Hyper-parameters. The hyper-parameters are determined through the grid search on the validation set. We use the ADAM optimization algorithm to train all the models. The learning rate for the condensation process is determined through a search over the set {0.01, 0.001, 0.0001}. The weight decay is 5e-4. Dropout is searched from [0, 1).
Computing Infrastructure. The codes are written in Python 3.9 and Pytorch 1.12.1. The operating system is Ubuntu 18.0, and all models are trained on GPUs. All experiments are conducted on a server with Intel(R) Xeon(R) CPUs (Gold 6128 @ 3.40GHz) and NVIDIA GeForce RTX 2080 Ti 11GB GPUs.
6.3 Experimental Results
Effectiveness. The condensed graphs generated by GC methods are evaluated to train the same 2-layer GCN [94] for node classification and the test accuracies with standard deviation are reported in Table IV. In the table, “Original graph” refers to the GCN performance which is trained on the original graph and we make the following observations. Firstly, GC methods are not sensitive to the compression rate and a larger condensed graph size does not strictly indicate better performance. Even with extremely small compression rates, GC methods can achieve promising performance. When comparing different optimization strategies, SFGC consistently outperforms other methods across all datasets, primarily due to its long-range trajectory matching. While GC-SNTK utilizes KRR to achieve competitive performance, the graph kernel in KRR is memory-intensive and does not scale well with large-scale graphs.
Generalization. We adhere to the settings commonly used in most GC studies and assess the generalization of GC methods across various GNN models for the node classification task. The GNN models evaluated include GCN [94], SGC [16], SAGE [9], APPNP [96], Cheby [97] and GAT [98]. The detailed accuracies are shown in Table V. Across various GNN architectures, most GC methods (i.e., GCond, GCDM, and SFGC) consistently achieve the best performance on GCN and SGC. This superior performance is attributed to the same convolution kernels [25] in the relay models of these GC methods, which share the same spectral characteristics [64] as those in GCN and SGC. Moreover, SFGC demonstrates a significant improvement over other methods, underscoring the effectiveness of trajectory matching in capturing essential information from the original graph and its superior generalization capabilities.
Efficiency. We report the condensation times of various GC methods on different datasets in Table VI. GCond employs gradient matching to align the original and condensed graphs, necessitating additional back-propagation steps for gradient computation relative to model parameters. In contrast, GCDM modifies the optimization objective and adopts distribution matching, thereby eliminating the need for gradient calculations and achieving faster condensation. Although GC-SNTK utilizes KRR to reduce iterative optimization in the inner loop and achieves rapid condensation on smaller datasets, it struggles to scale to larger graphs. SFGC requires training multiple GNNs on the original graph to provide the trajectories and necessitates multiple updates of the relay model to resemble the training trajectories. All these procedures are computationally intensive and result in extended condensation times.
TABLE VI: The condensation time (seconds) comparison of different GC methods. “OOM” denotes out-of-memory. $r$ is set as 2.60%, 1.80%, 0.25%, 0.50% and 0.10% for five datasets, respectively.
Dataset GCond GCDM GC-SNTK SFGC
Cora 498.9 27.3 30.8 2,524.4
Citeseer 479.3 62.0 19.4 3,877.0
Ogbn-arxiv 14,876.0 690.0 9,353.6 86,288.5
Flickr 1,338.3 209.5 562.6 52,513.0
Reddit 21,426.5 921.9 OOM 246,997.1
Fairness. Following FairGNN [91] and FGD [49], we evaluate GC methods on the Pokec-n dataset to assess group fairness issues within a binary sensitive feature setting. As shown in Table VII, GC methods amplify the graph data bias to varying degrees. GCond and GCDM not only achieve the lowest node classification accuracy but also exhibit the most pronounced bias. More advanced optimization strategies can effectively reduce this bias while preserving predictive performance.
TABLE VII: The comparison of accuracy (ACC, %), AUC (%), demographic parity ($\Delta_{DP}$, %) and equal opportunity ($\Delta_{EO}$, %) for various GC methods.
Metric GCond GCDM GC-SNTK SFGC
Original
graph
ACC $\uparrow$ 51.2±2.1 53.3±1.4 54.0±1.0 54.1±1.1 67.8±0.2
AUC $\uparrow$ 52.2±1.9 53.0±1.6 54.4±1.1 55.5±1.6 72.9±0.2
$\Delta_{DP}\downarrow$ 3.01±1.2 3.3±1.4 2.0±1.7 2.1±1.2 0.2±0.1
$\Delta_{EO}\downarrow$ 3.20±1.0 2.7±1.5 2.6±1.6 2.2±1.1 1.4±1.6
Robustness. Following the RobGC [58], we assess the resilience of GC methods to both random noise and adversarial attacks in the graph structure under the inductive setting. For random noise, this involves adding fake edges and removing real edges from the graph at random. Global adversarial attacks on the structure of the original graph are executed using PR-BCD [99]. Each method’s performance is evaluated at varying noise levels, defined by the ratio of changed edges in the original graph [100]. We assume uniform noise levels across real-world systems, meaning both training and test graphs are subjected to identical noise levels. As illustrated in Figure 10, “Original graph” denotes GNNs trained on the noisy training graph and tested on the noisy test graph. For the GC methods, GNNs are trained on the condensed graphs derived from the noisy training graphs. The adversarial attacks significantly impact classification performance more than random noise at equivalent noise levels. Among the GC methods, GCDM exhibits the lowest accuracy across different noise levels, whereas SFGC demonstrates the most robust performance against both random noise and adversarial attacks. However, as noise levels increase, all performances decline, and the gap between “Original graph” and the various GC methods increases, indicating that noise in the graph structure significantly degrades the quality of the condensed graphs produced.
Figure 10: The performance of GC methods under conditions of random noise and adversarial attacks on Ogbn-arxiv. “Original graph” indicates that GNNs are trained directly on the noisy training graph. In contrast, other methods involve training GNNs on the condensed graph. All GNNs are subsequently evaluated on the noisy test graph.
7 Applications and Resources
@@
Due to the efficiency in training GNN models and the capacity to conserve storage costs, GC contains vast application prospects and is increasingly being utilized across a variety of fields. In this section, we first enumerate the diverse applications of GC, including hyperparameter/neural architecture search, graph continual learning, federated learning, inference acceleration, recommender systems, and heterogeneous graphs. Consequently, this section details the open-source libraries for GC methods, facilitating the advancement of future research in this domain.
7.1 Practical Applications
7.1.1 Hyper-Parameter/Neural Architecture Search
@@
The hyper-parameter/neural architecture search [101] involves systematic experimentation with a range of hyper-parameters or architectures to identify the optimal combination for downstream tasks. This process typically requires the training of multiple models with varying hyper-parameters/architectures on the same dataset, incurring substantial computational costs. To mitigate the computational burden, HCDC [10] introduces GC in this problem and additionally generates the synthetic validation data by matching the hyper-parameter gradients. This ensures that the performance ranking of hyper-parameters/architectures on condensed datasets is consistent with that on the original datasets. By integrating GC, HCDC significantly reduces the time required for the search process.
7.1.2 Graph Continual Learning
@@
Graph continual learning [102] is the approach that focuses on the continuous acquisition of knowledge from dynamic and evolving graph data. In this realm, a critical challenge is the catastrophic forgetting problem, where a model’s performance on previously learned data deteriorates significantly when trained on new data. To mitigate this issue, CaT [27] and PUMA [65] incorporate GC into the process of graph continual learning, leveraging the condensed graphs to preserve historical information within a memory bank. To maintain a balance between historical and newly incoming data, CaT strategically trains the model exclusively on the condensed graph. This method not only addresses the challenge of data imbalance in graph continual learning but also significantly enhances learning efficiency.
7.1.3 Federated Learning